Job Costs Using a Plantwide Overhead Rate

Naranjo Company designs industrial prototypes for outside companies. Budgeted overhead for the year was $390,000, and budgeted direct labor hours were 26,000. The average wage rate for direct labor is expected to be $30 per hour. During June, Naranjo Company worked on four jobs. Data relating to these four jobs follow:

Job 39 Job 40 Job 41 Job 42
Beginning balance $23,300 $34,800 $19,700 $500
Materials requisitioned 19,800 20,400 10,600 12,600
Direct labor cost 10,900 17,500 5,250 3,500

Overhead is assigned as a percentage of direct labor cost. During June, Jobs 39 and 40 were completed; Job 39 was sold at 115 percent of cost. (Naranjo had originally developed Job 40 to order for a customer; however, that customer was near bankruptcy and the chance of Naranjo being paid was growing dimmer. Naranjo decided to hold Job 40 in inventory while the customer worked out its financial difficulties. Job 40 is the only job in Finished Goods Inventory.) Jobs 41 and 42 remain unfinished at the end of the month.

Required:

1. Set up a simple job-order cost sheet for all jobs in process during June.

Naranjo Company
Job-Order Cost Sheets
Job 39 Job 40 Job 41 Job 42
Balance, June 1 $ $ $ $
Beginning balance, Work-In-Process
Cost of goods sold
Ending balance, Work-in-Process
Total $ $ $ $

2. What if the expected direct labor rate at the beginning of the year was $24 instead of $30? What would the overhead rate be? If required, round your overhead rate answer to one decimal place.

New budgeted direct labor cost = $_____

New overhead rate = ______ % of direct labor cost

How would the cost of the jobs be affected?
Answer #1

1)Total Direct Labor cost = Direct labor hours * Rate per hour

                      = 26000 *30

                     = 780000

Overhead Rate =Estimated overhead /estimated direct labor cost

            = 390000 / 780000

           = .50 or 50% of direct labor cost

Naranjo Company

Job-Order Cost Sheets

Job 39 Job 40 Job 41 Job 42
Balance, June 1 23300 34800 19700 500
Material requisitioned 19800 20400 10600 12600
Direct labor cost 10900 17500 5250 3500
Overhead applied 10900*50%= 5450 17500*50%= 8750 5250*50%=2625 3500*50%= 1750
Total 59450 81450 38175 18350

2)

New budgeted direct labor cost = 26000 * 24= 624000

New Overhead rate = 390000/ 624000

                   = 62.5 % of Direct labor cost

With increase in overhead application rate from 50% to 62.5% ,Overhead applied will get increase as a result Total cost of job will increase.